OpenStudy (anonymous):

determine whether the series converges or diverges:

OpenStudy (anonymous):

\[\sum_{1}^{\infty} (lnk)/(e ^{\sqrt{k}})\]

OpenStudy (earthcitizen):

0

OpenStudy (anonymous):

what test did you use?

OpenStudy (earthcitizen):

when u find the limiting value \[U_{k+1}/U_{n}\]

OpenStudy (earthcitizen):

\[\left| U _{k+1}/U_{k} \right|\]

OpenStudy (anonymous):

can you kinda show me how you got zero?

OpenStudy (earthcitizen):

\[ U_{k} = \ln(k)/e^{k} , U_{k+1} = \ln(k+1)/e^{k+1} \]\[ \therefore \left| U_{k+1}/U_{k} \right| = \ln(k+1)/e^{k+1} \times e^{k}/\ln(k)\]

OpenStudy (earthcitizen):

\[ \ln(1) \times e^{\sqrt(k)-\sqrt(k+1)} = 0\]

OpenStudy (earthcitizen):

did that help ?

OpenStudy (anonymous):

great! thanks :)

OpenStudy (earthcitizen):

what was the answer in the text book ?

OpenStudy (anonymous):

it wasn't from my textbook so i don't know the right answer but that looks right.